OpenStudy (anonymous):

solve the system using any method: 6x-10y=12 -15x+25y=-30

OpenStudy (jdoe0001):

so, pick a method and use that :)

OpenStudy (anonymous):

how about the observe method?

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

i like elimination... especially in this problem

OpenStudy (jdoe0001):

yes, elimination would be simpler

OpenStudy (anonymous):

eliminate x

OpenStudy (anonymous):

just a suggestion...

OpenStudy (wolf1728):

A) 6x-10y=12 B) -15x+25y=-30 What about multiplying equation A by 2.5?

OpenStudy (anonymous):

yuck... mult 1st by 5 and 2nd by 2

OpenStudy (anonymous):

ok thank you both!

OpenStudy (wolf1728):

pgpilot326 - Isn't it better and easier just to multiply one equation?

OpenStudy (wolf1728):

Multiplying equation A makes it become A) 15x-25y=30 Then add this to equation B B) -15x+25y=-30

OpenStudy (anonymous):

just depends... in the end, the answer is the same.

OpenStudy (wolf1728):

Wow - it looks as though the 2 equations are identical - then no solution?

OpenStudy (anonymous):

not exactly... if they're the same then there are an infinite number of solutions

OpenStudy (wolf1728):

okay - yes - I consider it no solution - it's a lot less to write on a test or homework assignment. :-)

OpenStudy (anonymous):

but it's wrong to say no solution!

OpenStudy (wolf1728):

yes - you are right of course I would not want a professor or teacher to ask for ALL solutions - writer's cramp would ensue.

OpenStudy (anonymous):

you simply say all solutions are of the form (x, 3/5 x - 6/5) where x is any reall number.

OpenStudy (wolf1728):

okay - a mathematically elegant answer.